OpenStudy (raphaelll):

3t - 6 = -2t + 6 Solve for t

OpenStudy (anonymous):

first add +2t to both sides of the equation

OpenStudy (anonymous):

give me what you get

OpenStudy (raphaelll):

what do you mean?

OpenStudy (raphaelll):

3t - 6 + 2t = 2t +6+2t?

OpenStudy (anonymous):

almost, you forgot the negative on the 2t on the right hand side. Look at the original equation

OpenStudy (raphaelll):

ah my bad

OpenStudy (raphaelll):

what next?

OpenStudy (anonymous):

first of all, just remember what ever you do to one side you have to do the exact same to the other side

OpenStudy (anonymous):

first step move all the numbers that have a letter attached to them to any side second step you then want to move all numbers that don't have letters over to the opposite side from the numbers with letters the letters are called variables

OpenStudy (raphaelll):

so 3t+2t+-2t+2t=-6+6 is that right?

OpenStudy (anonymous):

yes and no, sorry i wasnt more specific ok lets go back to step number 1 so when you move something to the other side you move it by subtracting it or adding it, like -2x+7=x-5 lets say we want to move the x's to the right so we would start by adding 2x to both sides, the reason we add 2x is because we have a negative 2x on the left so if we add a positive 2x it gets rid of the negative 2x -2x+7=x-5 +2x = +2x 0+7 =x-5+2x 7=3x-5

OpenStudy (anonymous):

honey you'd add 2 to both side so that gives you 5t-6=6.then add 6 to each side, that will give you 5t=12 then divide by 5 on both side with will give you t=2.4

OpenStudy (anonymous):

@hisbaby this person needs help understanding how to do it, please don't confuse what I am attempting to teach

OpenStudy (anonymous):

teach him right and he might get it right.

OpenStudy (anonymous):

you're a woman of course you're going to argue

OpenStudy (raphaelll):

No need to fight guys, also i don't quite understand moving it to a side, could you explain how to do it using the actual problem?

OpenStudy (anonymous):

Check your messages raphaelll

OpenStudy (raphaelll):

replied :)

OpenStudy (anonymous):

3t - 6 = -2t + 6 +2t +2t what you do to one side you have to do to the other. see? when you add them, you get 5t-6=6 cause adding a +2 to a - 2 cancels it out. next you'd add 6 to both sides 5t-6=6 +6 +6 that makes the -6 cancel out and leaves you with 5t=12 next divide both sides by 5 5t/5=12/5 which gives you t=2.4

OpenStudy (anonymous):

ignore that gibberish

OpenStudy (accessdenied):

There are many ways to solve a problem, just as there are many ways to teach the solution. @eHoaX You need not dismiss that solution made by another user. Your methods are identical, only presented differently. Calling the method written 'gibberish' is not productive to either of your solutions. That sexist comment is equally unnecessary. @hisbaby You should respect the aforementioned user trying to help the student here, regardless of his/her approach. There are many ways to teach a student; not every person learns exactly the same way. So rather than dismiss their work entirely as "wrong", why not try and work together? Or at the very least, offer assistance rather than forcefully take over a question.